OpenStudy (anonymous):

Hi! Express the angle in terms of degrees, minutes, and seconds, to the nearest second. 14.878°

Directrix (directrix):

See if this makes sense. I am still processing it. Here's How: 1.The whole units of degrees will remain the same (i.e. in 121.135° longitude, start with 121°). 2.Multiply the decimal by 60 (i.e. .135 * 60 = 8.1). 3.The whole number becomes the minutes (8'). 4.Take the remaining decimal and multiply by 60. (i.e. .1 * 60 = 6). 5.The resulting number becomes the seconds (6"). Seconds can remain as a decimal. 1.Take your three sets of numbers and put them together, using the symbols for degrees (°), minutes (‘), and seconds (") http://geography.about.com/library/howto/htdegrees.htm

Directrix (directrix):

I got the 14 whole degrees, then 52', then 40.8". See what you get. @abigailmurcia

OpenStudy (debbieg):

Just think of it as a units conversion. There are 60 minutes in 1 degree, 60 seconds in 1 minute. So using the conversion factor \(\large \dfrac {60~min}{1~deg}\): \(\large 0.878~\cancel{deg}\times \dfrac {60~min}{1~\cancel{deg}}=52.68~min\) So now you have converted the "partial degrees" into minutes and partial minutes. Now you do the same thing with the conversion factor \(\large \dfrac {60~sec}{1~min}\) to convert the partial minutes to seconds.